2.Recording Direct Labor and Factory Overhead Chamlee Industries Inc. manufactures recreational vehicles. Chamlee Industries uses a job order cost system. The time tickets from May jobs are summarized below. Job 5-100 $8,320 Job 5-101 5,070 Job 5-102 6,240 Job 5-103 3,770 Factory supervision 2,275 Factory overhead is applied to jobs on the basis of a predetermined overhead rate of $35 per direct labor hour. The direct labor rate is $35 per hour. a. Determine the total factory labor costs transferred to Work in Process and Factory Overhead for May. If required, round your answers to nearest dollar
Answer #1

Solution:

A.)
Job-5-100 = $8,320
Job-5-100 = $5,070
Job-5-100 = $6,240
Job-5-100 = $3,770
Total = $23,400
Factory labor costs transferred to Work in Process = $23,400
Factory labor costs transferred to Factory Overhead = $2,275